### v1

@@ 7 Related Work
The widespread adoption of PTMs has created a complex software ecosystem, which researchers are beginning to analyze as a PTM supply chain (Jiang et al., 2022; Stalnaker et al., 2025), starting from upstream model hubs and packages and extending to their adaptation and integration within downstream software projects. While our work focuses on the downstream side of this chain, we review both perspectives to situate our contribution. We acknowledged some prior work in §2, and revisit relevant lines of research here, focusing first on hub-oriented (“upstream”) perspectives before turning to open-source ML-enabled systems (“downstream”) studies.
@@ 7.1 Model Hubs and Upstream Perspectives
However, these works primarily focus on upstream packages and hubs and do not investigate downstream source code integration, adaptation, or practical usage within real software projects. Our work addresses this gap by moving beyond hub-centric analysis to empirically characterize downstream PTM reuse patterns, examining how models are integrated, adapted, and interacted with in actual software systems.
@@ 7.2 Model Integration in ML-Enabled Systems
Recent research has analyzed integration and usage trends of ML models in open-source systems, often considering all ML models rather than focusing on PTMs (Sens et al., 2024; Nahar et al., 2025). For instance, Sens et al. (Sens et al., 2024) analyzed 2,928 ML-enabled systems on GitHub, finding that ML is primarily used for technical tasks rather than end-user functionality. Over half of the projects were ML libraries, with only 18% targeting end users. Extensive model reuse was observed, often via code duplication across repositories, introducing dependency challenges. However, the manual qualitative analysis was limited to 26 sampled systems, restricting the generalizability of insights into model reuse and model interactions. Importantly, the dataset was not curated specifically for PTM reuse. Another study curated 262 open-source ML products and analyzed 30 representative cases (Nahar et al., 2025), revealing distinctive architectural patterns, limited involvement of data scientists, and minimal adherence to best practices such as testing, monitoring, and pipeline automation. While these findings provide insights into production-oriented ML products, they do not focus on PTMs or cross-stage model reuse. Our distinct focus allowed us to examine PTM-specific reuse, addressing gaps in prior work.
@@
A growing line of research has examined the reuse of PTM specifically in downstream software projects (Jiang et al., 2024; Duan et al., 2024; Banyongrakkul et al., 2025). For instance, a recent work has examined PTM reuse specifically in downstream software projects by analyzing GitHub issues and pull requests (Banyongrakkul et al., 2025). The taxonomies derived from this study identified challenges in model usage, adaptation, configuration, and resource management, showing that PTM-related issues take longer to resolve than non-PTM issues. While these studies highlight barriers, they do not capture systematic PTM integration within software systems. Other research has documented metadata for 281,638 PTM packages and 28,575 downstream repositories, including license information, revealing that 0.24% of licenses between PTMs and their dependent projects are inconsistent (Jiang et al., 2024). Tools like ModelGo complement this by assessing license conflicts, improper license choices, and obligations across software, datasets, and PTMs (Duan et al., 2024). These findings highlight the gaps in understanding downstream adoption and reuse of PTM that we address in this work.
@@ 8 Conclusion
This study provides the first large-scale empirical characterization of pre-trained models (PTMs) as a new form of software dependency, which we term Software Dependencies 2.0. By analyzing a representative sample of 401 GitHub projects, we examined how PTMs are managed and integrated into real-world software. Our results show that PTM dependencies are often inconsistently specified across code, documentation, and configuration files, and adaptation is common, ranging from as-is reuse to architectural modification. We identified three dominant PTM reuse pipeline types—feature extraction, generative, and discriminative—and observed frequent interactions among PTMs and with other models, introducing significant architectural complexity.
These findings highlight the technical challenges of managing Software Dependencies 2.0, including maintainability, reproducibility, and system reliability. By conceptualizing PTMs as first-class, modular components, our work informs the development of improved tools, standards, and best practices to support reliable integration and long-term maintenance. Future research should investigate automated support for PTM dependency management, examine reuse and adaptation in industrial-scale ML systems, and further explore multi-model interactions to advance understanding and engineering of Software Dependencies 2.0.
9 Data Availability
Our artifact is publicly available at https://github.com/RISElabQueens/PTMReuseInOSS. It contains the dataset of PTM usage with annotated labels and scripts for analysis, figures and plots.
Acknowledgement
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) [funding reference number: RGPIN-2019-05071] and a gift from Socket, Inc.

### v2

@@ 7 Related Work
The widespread adoption of PTMs has created a complex software ecosystem, which researchers are beginning to analyze as a PTM supply chain (Jiang et al., 2022; Stalnaker et al., 2025), starting from upstream model hubs and packages and extending to their adaptation and integration within downstream software projects. While our work focuses on the downstream side of this chain, we review both perspectives to situate our contribution. We acknowledged some prior work in §2, §5 and revisit relevant lines of research here, focusing first on hub-oriented (“upstream”) perspectives before turning to open-source ML-enabled systems (“downstream”) studies.
@@ 7.1 Model Hubs and Upstream Perspectives
These works primarily focus on upstream packages and hubs and do not investigate downstream source code integration, adaptation, or practical usage within real software projects. Our work addresses this gap by moving beyond hub-centric analysis to empirically characterize downstream PTM reuse patterns, examining how models are integrated, adapted, and interacted with in actual software systems.
@@ 7.2 Model Integration in ML-Enabled Systems
Recent research has analyzed integration and usage trends of ML models in open-source systems, often considering all ML models rather than focusing on PTMs (Sens et al., 2024; Nahar et al., 2025). For instance, Sens et al. (2024) analyzed 2,928 ML-enabled systems on GitHub, finding that ML is primarily used for technical tasks rather than end-user functionality. Over half of the projects were ML libraries, with only 18% targeting end users. Extensive model reuse was observed, often via code duplication across repositories, introducing dependency challenges. However, the manual qualitative analysis was limited to 26 sampled systems, restricting the generalizability of insights into model reuse and model interactions. Importantly, the dataset was not curated specifically for PTM reuse. Meanwhile, Nahar et al. (2025) study curated 262 open-source ML products and analyzed 30 representative cases, revealing distinctive architectural patterns, limited involvement of data scientists, and minimal adherence to best practices such as testing, monitoring, and pipeline automation. While these findings provide insights into production-oriented ML products, they do not focus on PTMs or cross-stage model reuse. Our distinct focus allowed us to examine PTM-specific reuse, addressing gaps in prior work.
Prior work has examined the broader landscape of ML library usage and software engineering challenges in ML-enabled systems. Dilhara et al. (2021) studied ML library usage and evolution, analyzing how developers interact with ML libraries and how these libraries evolve over time. In the context of what Karpathy (2017) termed “Software 2.0,” their work provides insights into library-level dependencies and API usage patterns. While their focus is on ML libraries broadly, our work specifically targets PTMs as a distinct form of dependency, examining not just library usage but the full integration pipeline, including model-specific stages (e.g., fine-tuning, inference) and cross-stage interactions between PTMs and other components. More recently, Shivashankar et al. (2025) surveyed maintainability and scalability challenges in ML systems, identifying pain points and solutions. While their work provides a comprehensive overview of ML system challenges, our study contributes empirical evidence specifically for PTM dependencies—characterizing how these dependencies are structured (RQ1), organized within reuse pipelines (RQ2), and composed with other models (RQ3). Our PTM-specific focus reveals patterns and maintenance challenges unique to model-centric dependencies that may not fully emerge in broader ML system studies.
@@
A growing line of research has examined the reuse of PTM specifically in downstream software projects (Jiang et al., 2024; Duan et al., 2024; Banyongrakkul et al., 2025). For instance, Banyongrakkul et al. (2025) examined PTM reuse specifically in downstream software projects by analyzing GitHub issues and pull requests. Their taxonomies recorded challenges in model usage, adaptation, configuration, and resource management, showing that PTM-related issues take longer to resolve than non-PTM issues. While these studies highlight barriers, they do not capture systematic PTM integration within software systems. Other research has documented metadata for 281,638 PTM packages and 28,575 downstream repositories, including license information, revealing that 0.24% of licenses between PTMs and their dependent projects are inconsistent (Jiang et al., 2024). Tools like ModelGo complement this by assessing license conflicts, improper license choices, and obligations across software, datasets, and PTMs (Duan et al., 2024). These findings highlight the gaps in understanding downstream adoption and reuse of PTM that we work to address in this work.
@@ 8 Conclusion
This study provides the first large-scale empirical characterization of pre-trained models (PTMs) as a new form of software dependency, which we term Software Dependencies 2.0. By analyzing a representative sample of 401 GitHub projects, we examined how PTMs are managed and integrated into real-world software. Our results show that PTM dependencies are often inconsistently specified across code, documentation, and configuration files, and adaptation is common, ranging from as-is reuse to architectural modification. We identified three dominant PTM reuse pipeline types — feature extraction, generative, and discriminative — and observed frequent interactions among PTMs and with other models, introducing significant architectural complexity beyond conventional library reuse. Taken together, these findings provide empirical grounding for Software Dependencies 2.0 as model-centric, pipeline-embedded, and interaction-driven dependencies.
Our results suggest that managing Software Dependencies 2.0 requires practices that extend beyond traditional dependency management, including improved traceability of model artifacts, explicit versioning, and support for multi-model coordination. By conceptualizing PTMs as first-class, modular components rather than opaque add-ons, the software engineering community can develop tools, standards, and architectural guidance tailored to model-centric systems. More broadly, advancing the theory and engineering of Software Dependencies 2.0 will be critical to support reliable integration and long-term maintenance as model-centric components become increasingly central to modern software systems. Our results guide future research: to investigate automated support for PTM dependency management, examine reuse and adaptation in industrial-scale ML systems, and further explore multi-model interactions to advance understanding and engineering of Software Dependencies 2.0.